OpenStudy (anonymous):

Transform each polar equation an equation in rectangular coordinates identify its shape. 1) rsin(theta)=3 (I just need the shape) 2) r = 2cos(theta) (I just need the coordinates) 3) r = (4/(2cos(theta) - 3sin(theta))) (I just need the coordinates) 4) rcsc(theta) = 5 (I just need the shape)

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

I can't use the draw function !

OpenStudy (anonymous):

Ok, then can you help me with the coordinates then?

OpenStudy (anonymous):

What do you mean by "needing the coordinates"? You have to convert from one coordinate \(\textbf{system}\) to another.

OpenStudy (anonymous):

Exactly...

OpenStudy (anonymous):

Doesn't the rectangular coordinate system look like regular coordinates? Ex: (x, y)

OpenStudy (anonymous):

Yes, rectangular refers to the coordinate grid. The polar coordinate system describes the same points in a different way.

OpenStudy (anonymous):

Specifically, points in the rectangular coordinate system taking the form \((x,y)\) are expressed in the form \((r,\theta)\) in polar coordinates, where \(x=r\cos\theta\) and \(y=r\sin \theta\).

OpenStudy (anonymous):

I thought y was theta not rsin(theta)

OpenStudy (anonymous):